OpenStudy (anonymous):

f(x)=x^2+3x-5 Solve (f(x+h)-f(x))/h h cannot = 0

OpenStudy (anonymous):

bunch of algebra ready?

OpenStudy (anonymous):

let's do this

OpenStudy (anonymous):

ok first we need \[f(x+h)\] and since \[f(x)=x^2+3x+5\] then \[f(x+h)=(x+h)^2+3(x+h)+5\] now comes the first part of algebra

OpenStudy (anonymous):

square carefully and get \[f(x)=x^2+2xh+h^2+3x+3x+5\]

OpenStudy (anonymous):

now we need \[f(x+h)-f(x)=x^2+2xh+h^2+3x+3h+5-(x^2+3x+5)\]

OpenStudy (anonymous):

sure looks like a derivative, df/dx = 2x + 3, etc.

OpenStudy (anonymous):

don't forget the parentheses about the last part distribute the minus sign and get \[x^2+2xh+h^2+3x+3h+5-x^2-3x-5\] combine like terms now

OpenStudy (anonymous):

@douglaswinslowcooper in particular, it's a difference quotient; I assume that if they're just working with these, they may not know the concept of a limit yet, and hence you can't do derivatives.

OpenStudy (anonymous):

everything without an \(h\) will go and you get \[f(x+h)-f(x)=2xh+3h+h^2\]

OpenStudy (anonymous):

finally divide by \(h\) to ge t \[\frac{f(x+h)-f(x)}{h}=\frac{2xh+3h+h^2}{h}\] divide each term by \(h\) leaving \[2x+3+h\]

OpenStudy (anonymous):

like a said, bunch of algebra, and there is lots of room for mistake here especially squaring \((x+h)^2\) and also making sure to put \(-f(x)\) in parentheses